### v1

# Convex Markov Games: A Framework for Fairness, Imitation, and Creativity in Multi-Agent Learning

Abstract
Expert imitation, behavioral diversity, and fairness preferences give rise to preferences in sequential decision making domains that do not decompose additively across time. We introduce the class of convex Markov games that allow general convex preferences over occupancy measures. Despite infinite time horizon and strictly higher generality than Markov games, pure strategy Nash equilibria exist under strict convexity. Furthermore, equilibria can be approximated efficiently by performing gradient descent on an upper bound of exploitability. Our experiments imitate human choices in ultimatum games, reveal novel solutions to the repeated prisoner’s dilemma, and find fair solutions in a repeated asymmetric coordination game. In the prisoner’s dilemma, our algorithm finds a policy profile that deviates from observed human play only slightly, yet achieves higher per-player utility while also being three orders of magnitude less exploitable.
1 Introduction
It has been a characteristic assumption in multi-agent reinforcement learning that agent rewards are a discounted sum of per-period rewards (Littman, 1994). This article proposes a framework beyond such decomposed reward, and argues its algorithmic tractability and utility.
Already early game theory has observed that maximizing decomposable rewards may lead to a high number of equilibria, of very different desirability for a designer, (Friedman, 1971). In addition, many search problems have other desiderata, e.g., diversity of play, that cannot be expressed as reward maximization (Zahavy et al., 2023). In addition, such utilities do not permit the designers to find approximate equilibria that have desirable features such as being human-like (imitation), or that are creative in that their trajectories visit many parts of the space, or to include fairness concerns.
In this paper, we define a general yet tractable framework for the inclusion of fairness, imitation and creativity components in multi-agent learning.
A crucial object for our analysis is the occupancy measure $\mu$. Occupancy measures are the long-term distributions of states and actions in different time periods, taking into account discounting. That is, the weight of a state-action profile $(s,(a_{1},a_{2},\dots,a_{n}))$ is obtained by summing the probabilities that $s_{t}=s$ and $a_{1t}=a_{1},a_{2t}=a_{2},\dots,a_{nt}=a_{n}$ discounted by $\gamma^{t}$. In general, the class of utilities that we consider here are concave functions of occupancy measures (equivalently, players’ “losses” or negated utilities are convex). Restricting to the class of linear utility functions, we recover Markov games, that is, the classic model of multi-agent reinforcement learning (Littman, 1994).
We describe our more general framework in the context of three classes of utility functions. For this discussion, we consider a linear reward function $r\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$, which we view as the reward from a game. 1) A first class is those utilities that favour dispersion of the state-action profiles, $u(\mu)=r^{\top}\mu+\tau H(\mu)$, where $H$ is an entropy measure such as Shannon, Rényi, or Tsallis entropy, and $\tau>0$. Whereas entropy of player policies $H(\pi)$ measures randomness of their action selection in each state and only weakly correlates with diversity of state visitation, entropy of player occupancy measures directly measures both, especially extending to direct measurement of the latter. 2) Surprisingly, in addition to giving a bonus for exploring, an apparently opposing utility can be accommodated while maintaining tractability, as we show: imitation utilities, $u(\mu)=r^{\top}\mu-\tau d(\mu,\nu)$ for some target occupancy measure $\nu$ and some distance measure $d$ such as reverse Kullback-Leibler divergence, Wasserstein, or total variation metric. This allows for equilibria that are close to others. 3) Finally, fairness can be encouraged through penalty functions, e.g., $u(\mu)=r^{\top}\mu-\tau\mu^{\top}\Sigma\mu$ for some $\Sigma\succeq 0$. We explore each of the three classes in our experiments in Section 5.
Linear Loss Convex Loss (Concave Utility)
Single-Agent Markov Decision Process (MDP) convex MDP
Multi-Agent Markov (Stochastic) Game (Ours) convex Markov Game
Table 1: We introduce the convex Markov Game (cMG) to fill a gap at the intersection of multi-agent and reinforcement learning research. We demonstrate later how to use cMGs to model three classes of utilities for a linear reward function $r$: (a) Creativity, using an entropy bonus term, (b) Imitation, using a distance penalty to a reference occupancy measure, and (c) Fairness, using a positive definite penalty. This paper establishes existence of equilibria for such utilities, proposes a differentiable loss for its computation, and demonstrates the performance of this loss in the presence of imitation, creativity, and fairness desiderata.
We initiate our study of convex Markov games with three contributions:
  • $\bullet$
    First, we show that under strictly concave utilities, pure-strategy Nash equilibria exist. This is the foundation for training policies in multi-agent reinforcement learning that do not require randomizing over multiple models.
  • $\bullet$
    Second, we provide a loss that bounds the deviation from equilibrium in a convex Markov game. Minimizing this loss allows for the computation of equilibria.
  • $\bullet$
    Finally, we show that convex Markov games are capable of generating creative, human-like and fair equilibria in repeated games.
The rest of the article is structured as follows. We define the class of convex Markov Games in section 2. We show that pure-strategy Nash equilibria exist in section 3. Our analysis of a loss for equilibrium computation is in section 4, which we put to practice in our experiments in section 5. We discuss related literature in section 6, and conclude in section 7.
2 Convex Markov Games
The main definition of this article is the convex Markov Game (cMG).
Definition 1.
A convex Markov game is given by a $6$-tuple $\mathcal{G}=\langle\mathcal{S},\mathcal{A}=\operatorname*{\vartimes}_{i=1}^{n}\mathcal{A}_{i},P,u,\gamma,\mu_{0}\rangle$:
  • $\bullet$
    Players $i=1,2,\dots,n$,
  • $\bullet$
    a finite state space $\mathcal{S}$, and an initial state distribution $\mu_{0}\in\Delta^{\mathcal{S}}$,
  • $\bullet$
    finite action spaces $\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{n}$,
  • $\bullet$
    a state transition function $P\colon\mathcal{S}\times(\operatorname*{\vartimes}_{i=1}^{n}\mathcal{A}_{i})\to\Delta^{\mathcal{S}}$,
  • $\bullet$
    a discount factor $\gamma\in[0,1)$, and
  • $\bullet$
    a set of continuous utilities with each $u_{i}$ concave in the $i$th player’s occupancy measure, $u_{i}\colon(\operatorname*{\vartimes}_{j=1}^{n}\Delta^{\mathcal{S}\times\mathcal{A}_{j}})\to\mathbb{R}$.
The Policy View
Players $i=1,\ldots,n$ choose policies $\pi_{i}\colon\mathcal{S}\times\mathcal{A}_{i}\to[0,1]$. A policy profile $\pi=(\pi_{1},\pi_{2},\dots,\pi_{n})$ induces a sequence of states and joint actions $(s_{t})_{t\in\mathbb{N}}$ and $(a_{t})_{t\in\mathbb{N}}$, and a state-action occupancy measure
$\mu^{\pi}(s,a)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}(s_{t}=s,a_{t}=a|\mu_{0},\pi,P).$
We can recover state-action occupancies from a matrix equation. Let $P^{\pi}$ be the state transition matrix under $\pi$. Then the state occupancy measure $\mu^{s}(s)=\sum_{a}\mu(s,a)$ can be written as a function of $\pi$ in matrix notation as
$\mu^{s}(\pi)$ $=(1-\gamma)\Big([I-\gamma P^{\pi}]^{-1}\mu_{0}\Big),$ (1)
and the state-action occupancy can be recovered as $\mu(s,a)=\mu^{s}(s)\cdot\pi(a|s)$. The marginal on this probability for only agent $i$’s actions can be recovered as
$\mu_{i}(\pi_{i},\pi_{-i})$ $=(1-\gamma)([I-\gamma P^{\pi}]^{-1}\mu_{0}\mathbf{1}_{|\mathcal{A}_{i}|}^{\top})\odot\pi_{i},$ (2)
where $\odot$ denotes the Hadamard product and $\pi_{-i}$ indicates all policies except player $i$’s.
Players may choose to randomize over stochastic policies. We say that a player’s strategy $\rho_{i}$ is a mixed-strategy if it is a distribution over policies. For example, let $\pi_{i}^{(a)}$ and $\pi_{i}^{(b)}$ both be stochastic policies. Then an example of a mixed-strategy is one in which player $i$ plays $\pi_{i}^{(a)}$ and $\pi_{i}^{(b)}$ with equal probability $\nicefrac{{1}}{{2}}$. If all probability mass of $\rho_{i}$ is on a single policy $\pi_{i}$, we call $\rho_{i}$ a pure-strategy and write $\pi_{i}$ directly.
We say that a (random) policy profile $\rho=(\rho_{1},\rho_{2},\dots,\rho_{n})$ is a (mixed-strategy) Nash equilibrium if for all agents $i=1,2,\dots,n$,
$\mathbb{E}_{\pi\sim\rho}[u_{i}(\mu(\pi_{i},\pi_{-i}))]\geq\mathbb{E}_{\pi_{i}^{\prime}\sim\rho_{i}^{\prime},\pi_{-i}\sim\rho_{-i}}[u_{i}(\mu(\pi_{i}^{\prime},\pi_{-i}))],$
for any policy $\rho_{i}^{\prime}$. If $\rho$ is a pure strategy profile as defined above, we call it a pure-strategy Nash equilibrium.
We call any policy $\tilde{\pi}_{i}$ (or occupancy measure $\tilde{\mu}_{i}$) player $i$’s best response to a profile $\rho$ if it achieves an expected utility that is maximal among all policies (respectively, occupancy measures). Note best responses are the solutions to individual player’s MDP with all other player policies held fixed, and so they necessarily exist as pure strategies: $\tilde{\pi}=\operatorname*{arg\,max}_{\pi_{i}^{\prime}}\mathbb{E}_{\pi_{-i}\sim\rho_{-i}}[u_{i}(\mu(\pi_{i}^{\prime},\pi_{-i}))]$
The Occupancy Measure View
In our analysis and the algorithms proposed in this article, it will be helpful to frame and solve problems directly in the space of occupancy measures $\mathcal{U}$.
Given opponent policy profiles $\pi_{-i}$ we can define the probability kernel $P_{i}^{\pi_{-i}}$ as
$P_{i}^{\pi_{-i}}(s^{\prime}|s,a_{i})$ $=\sum_{a_{-i}\in\mathcal{A}_{-i}}P(s^{\prime}|s,a_{i},a_{-i})\prod_{j\neq i}\pi_{j}(a_{j}|s).$
Given these, we can reformulate an individual agent’s decision problem as
$\begin{split}\max_{\mu_{i}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}_{i}|}}&u_{i}(\mu_{i},\pi_{-i})\\ s.t.\quad\mu_{i}(s,a)&\geq 0\,\,\forall\,\,s\in\mathcal{S},a\in\mathcal{A}_{i}\\ \sum_{a_{i}\in\mathcal{A}_{i}}\big(I-\gamma P^{\pi_{-i}}_{i}(\cdot,a_{i},\cdot)\big)\mu_{i}(\cdot,a_{i})&=(1-\gamma)\mu_{0}(s)\,\,\forall\,\,s\in\mathcal{S}\end{split}$ (3)
where $P^{\pi_{-i}}_{i}(\cdot,a_{i},\cdot)$ is a next-state by state transition matrix and $\mu_{i}(\cdot,a_{i})$ is a vector denoting the probability of taking action $a_{i}$ in every state. Because $u_{i}$ are concave in $\mu_{i}$, and given the linearity of the constraints in (3), this problem is convex, motivating the name of convex Markov Games (cMGs). If $u_{i}(\mu_{i},\pi_{-i})=r^{\top}\mu_{i}$ as mentioned in the introduction, we recover the vanilla Markov game framework, however cMGs allow a much wider class of utilities (e.g., ones that include entropy of the occupancy measure). Notice that the feasible set in (3) is defined by constraints that depend on $\pi_{-i}$. For convenience, we define the so-called action correspondence, $\mathcal{U}_{i}=\mathcal{M}_{i}(\pi_{-i})$, which maps every profile of opponent policies to the feasible set of occupancy measures for player $i$, i.e, problem (3) can be succinctly written
$\max_{\mu_{i}\in\mathcal{M}_{i}(\pi_{-i})}u_{i}(\mu_{i},\pi_{-i}).$
We will use the feature that policies can be recovered from occupancy measures as
$\pi_{i}(\mu_{i})(a|s)$ $=\begin{cases}\frac{\mu_{i}(s,a)}{\sum_{a^{\prime}}\mu_{i}(s,a^{\prime})}&\text{ if }\sum_{a^{\prime}}\mu_{i}(s,a^{\prime})>0\\ \text{\emph{arbitrary}}&\text{ otherwise}.\end{cases}$ (4)
3 Existence of Nash Equilibrium
We first show that under fairly general assumptions, mixed Nash equilibria exist. We then argue that for the purposes of learning, pure-strategy Nash equilibria are particularly desirable, and show that these exist when utilities are strictly concave.
Proposition 1.
Mixed-strategy Nash equilibria exist in convex Markov Games.
This statement follows easily by using the policy view. Each player’s optimization problem in terms of policies is:
$\max_{\pi_{i}\in(\operatorname*{\vartimes}_{s=1}^{|\mathcal{S}|}\Delta^{\mathcal{A}_{i}})}$ $u_{i}(\mu_{i}(\pi_{i},\pi_{-i}),\pi_{-i}).$ (5)
Appendix B shows that all $u_{i}$ are continuous, differentiable functions of $(\pi_{i},\pi_{-i})$, and since the strategy sets are compact, the existence of mixed-strategy Nash equilibria follows from classical existence results (Glicksberg, 1952).
While existence of mixed-strategy equilibria are a descriptively helpful tool, they have limitations for applications in learning. In particular, learning a continuous distribution over stochastic policies would practically require function approximation or defining a mesh over the space, greatly increasing the complexity of the learning problem; this challenge has been broached but not sufficiently solved in the context of, for example, generative adversarial networks (Arora et al., 2017). Therefore, we provide another statement that ensures the existence of pure-strategy Nash equilibria.
Theorem 1.
Given a convex Markov game with unique best response correspondences in the space of occupancy measures, Nash equilibria exist in stationary stochastic policies (and hence state-action occupancy measures).
The proof of this statement relies on the continuity of players’ utilities and compactness of their strategy sets as before, but also the uniqueness of best responses. [^1] While our proof technique follows that of (Rosen, 1965) in relying on the Kakutani fixed-point theorem (Kakutani, 1941), the theory of concave games does not directly apply to our proposed cMG framework as agent utilities are concave in their occupancy measures, but not concave in their policies. The classes of strictly convex functions and invex functions (Hanson, 1981) are two that give rise to unique best response correspondences. Most definitions of (neg)entropy and distance metrics satisfy strict convexity. Any penalty function given by a quadratic with positive definite matrix $\Sigma$ does as well.
We briefly address the limitation of unique best response correspondences with a more detailed discussion in the Appendix G. While concave utilities give rise to convex best response correspondences in occupancy measure space, the image of these sets in policy space may be non-convex. The foundation of many equilibrium existence results rely on the assumption of a mapping from points (a strategy profile) to convex sets (a set of strategy profiles) via the Kakutani fixed point theorem. This core assumption is broken in the more general cMG setting. Proving existence of Nash equilibria in this more general setting is desired as the allowance of concave (not-strictly) utilities would enable the specification of “safe” occupancy measure sets (Miryoosefi et al., 2019).
4 Computation of Equilibria
While equilibria exist, they may be hard to compute with general solvers due to the size of the equilibrium computation problem and the expense of repeatedly solving convex programs. We present a gradient-based approach for equilibrium computation. We first use the occupancy measure view to derive a measure of utility-gain ($\epsilon_{i}$) from player best responses, and then reparameterize this loss to policies to circumvent challenges arising from non-convexity of the set of occupancy measures (Appendix G.2).
We provide an upper bound on exploitability ($\epsilon$) as a loss for equilibrium computation, defined as
$\epsilon$ $=\max_{i=1,\dots,n}\epsilon_{i}\qquad\text{where}\qquad\epsilon_{i}=\left(\max_{z\in\mathcal{M}_{i}(\mu_{-i})}u_{i}(z,\mu_{-i})\right)-u_{i}(\mu_{i},\mu_{-i}).$ (6)
Exploitability measures the most any player can gain by unilaterally deviating. Mechanistically, it corresponds to each player solving problem (3)—a constrained, convex optimization problem—and then reporting the value they attained beyond that of their strategy under the approximate equilibrium profile $(\mu_{1},\ldots,\mu_{n})$. In particular, an occupancy measure $(\mu_{1},\ldots,\mu_{n})$ with vanishing exploitability is a pure-strategy Nash equilibrium; the same NE profile from the policy view can be derived from equation 4.
The next result extends that of Gemp et al. (2024) from the normal-form game setting to show that exploitability is bounded from above by a constant depending on the action space size, and a projected gradient (compare white vector to one behind it in Figure 1). This bound is only tight for solutions in the interior of the simplex of occupancy measures. This is intuitive as boundary solutions may feature non-zero gradients in the tangent dimension despite having zero exploitability. The following result bounds the exploitability of a cMG with utilities $u_{i}$ using players’ utility-gradients of a cMG with a small amount of entropy regularization, i.e. $u_{i}^{\tau}(\mu_{i},\pi_{-i})=u_{i}(\mu_{i},\pi_{-i})+\tau H(\mu_{i})$, where $H$ denotes Shannon entropy.
(a) Occupancy Measure Space ($\mathcal{U}$)
(b) Policy Space ($\Pi$)
Figure 1: (A *not strictly* convex Markov Game) A 22-player, 22-state, 22-action convex Markov game with concave, but not strictly concave utilities (not shown). Colors help visualize the effect of the nonlinear transformation from occupancy space to policy space (see equation 4). See additional environment details in Appendix G.1. (1(a)) The feasible set of occupancy measures for player 11 given a fixed policy for player 22 with best response region is shown in black. Player $1$’s occupancy gradient $\nabla^{i}_{\mu_{i}}$ points off the 33-simplex so is not shown. Instead, we show $\nabla^{i}_{\mu_{i}}$ projected onto the tangent space of the feasible set in white, i.e., $\Pi_{T\mathcal{M}_{i}}(\nabla^{i}_{\mu_{i}})$. The vector shadowing the white one is $\nabla^{i}_{\mu_{i}}$ projected onto the tangent space of the simplex; note this vector points off the feasible set (into the page). (1(b)) A 22-d slice of player $1$’s feasible policy space. The set of best response policies is non-convex when viewed in policy space.
Theorem 2 (Low Temperature Approximate Equilibria are Approximate Nash Equilibria).
Assume that $\mu_{0}$ has full support. Let $\nabla^{i\tau}_{\mu_{i}}$ be player $i$’s entropy regularized gradient and $\pi$} be an approximate equilibrium of the entropy-regularized game. Then,
$\epsilon_{i}({\pi})\leq\tau\log(|\mathcal{A}_{i}|)+\sqrt{2}\|\Pi_{T\mathcal{M}_{i}}(\nabla^{i\tau}_{\mu_{i}})\|,$
where $\Pi_{T\mathcal{M}_{i}}$ is a projection matrix on the tangent space of $\mathcal{M}_{i}$, and $\nabla^{i\tau}_{\mu_{i}}$ is the gradient of $u_{i}^{\tau}$ with respect to $\mu_{i}$.
Inspired by this bound, we define the following projected-gradient loss (PGL) function for cMGs:
[…]
As a consequence of Theorem 2, we obtain that
$\epsilon({\pi})$ $\leq\tau\log(|\mathcal{S}|^{n}|\mathcal{A}|)+\sqrt{2n\mathcal{L}^{\tau}({\pi})}.$ (8)
Observe that this analysis purely leverages the occupancy measure view, that is, view the problem as parameterized by $\mu_{i}$. The space of feasible occupancy measures $\mu$, unfortunately, is a nonlinear manifold, and minimization of $\mathcal{L}^{\tau}$ may be challenging. While for any fixed $\pi_{-i}$ the contraints in equation 3 are linear, the joint constraints for feasible occupancy measures are not (Appendix G.2).
Instead of maximizing occupancy measures, we can consider optimizing directly over the space of policies $\pi_{i}$. Each player’s policy is subject to simplex constraints that are independent of the other players. $\mathcal{L}^{\tau}({\pi})=\mathcal{L}^{\tau}({\mu}({\pi}))$ with ${\mu}({\pi})$ defined in Section 2. In this way, we combine the two distinct strengths of the policy and occupancy measure views. In the policy view, player strategy sets (policy sets) are independent and convex, making for simple independent updates and projections back to the feasible sets. In the occupancy measure view, we enjoy convexity of player losses which enable derivation of upper bounds on exploitability.
Note that our proposed loss is composed of a projection operator, the gradients of our concave utilities $u_{i}$, and the mapping from policies to occupancy measures. We prove in several lemmas in the appendix that all of these components are differentiable assuming that the utilities are not only concave, but also differentiable [^2] Strict differentiability of utilities is not required for autodiff libraries, e.g., JaX (Bradbury et al., 2018)..
We can give more intuition for the analysis of the projection operator. If $A(\mu_{-i})\in\mathbb{R}^{\lvert\mathcal{S}\rvert\times(\lvert\mathcal{S}\rvert\cdot\lvert\mathcal{A}_{i}\rvert)}$ such that $A(\mu_{-i})\mu_{i}=(1-\gamma)\mu_{0}$ represents the linear equality constraints in equation 3, then the projection matrix is given by
$\Pi_{T\mathcal{M}_{i}(\mu_{-i})}$ $=I_{s\times s}-A^{\top}(AA^{\top})^{-1}A.$ (9)
For ease of notation, we omit the dependence of $A$ on $\mu_{-i}$. $\Pi_{T\mathcal{M}_{i}(\mu_{-i})}$ is differentiable if and only if $A(\mu_{-i})$ has full row rank (i.e., rank $|\mathcal{S}|$). The following result shows that $A(\mu_{-i})$ has full row rank.
Lemma 1.
For any $\gamma<1$, the equality constraint matrix in equation 3 has full row rank ($|\mathcal{S}|$).
This yields a differentiable loss and allows for optimization using automatic differentiation. If we represent agent policies in an unconstrained space with logits, we can then minimize $\mathcal{L}^{\tau}$ directly with respect to agent policies using our preferred unconstrained optimizer, e.g., Adam. We refer to this approach as PGL for projected gradient loss minimization.
5 Experiments
We test three variants of utilities that do not decompose linearly in four domains. We compare against four baseline algorithms, and compare the resulting exploitability and policy profiles.
Baselines.
We compare against four baselines. The first baseline, $\min\epsilon$, directly minimizes exploitability with respect to the agent policies using a differentiable convex optimization package cvxpylayers in JAX (Agrawal et al., 2019; Bradbury et al., 2018). In the second baseline, Sim, all players simultaneously run gradient descent on their losses with respect to their policies and we report the performance of the running average of the policy trajectory. Policies at each state are represented in $\mathbb{R}^{\lvert\mathcal{A}_{i}\rvert-1}$ as a softmax over $|\mathcal{A}_{i}|-1$ logits with the last logit fixed as 0. In the third, RR, agents alternate gradient descent steps in round-robin fashion. We also compare against the sgamesolver (Eibelshäuser & Poensgen, 2023) package of homotopy methods for Markov games.
Hyperparameters.
We use an exponentially decaying annealing schedule of $\tau=10^{-t}$ for $t\in[0,\ldots,7]$, with $1000$ iterations at each temperature $\tau$ for our entropy regularization. Adam’s internal state is not reset after annealing. We initialize to uniform policies (and whatever occupancy measure that induces) unless otherwise specified. All experiments except pathfinding were run on a single CPU and take about a minute to solve although exact exploitability reporting via cvxopt (Diamond & Boyd, 2016) increases runtime approximately $10\times$. Pathfinding was run on a single GPU.
Figure 2: (Approximate PGL Equilibrium) Arrows indicate each player’s most likely action (4 cardinal directions + “Stay”) with their color indicating their probability (lightest occurs at $0.20$ probability, darkest at $1.0$). Moves to the same location are awarded randomly to one of the agents. The goal state is marked by B and G where both green reaches the top right corner and blue the top left.
Figure 3: (Convergence to Nash equilibria) $R$ and PGL descent yield the lowest exploitability yet converge to different equilibria. PGL drops coincide with temperature annealing. $\min\epsilon$ crashes, marked by stars.
Domains.
We consider five domains: The first is a multiagent pathfinding problem in a grid world.The second domain is the classic two-player, iterated prisoner’s dilemma (compare Tucker & Straffin Jr (1983)) where agents may choose to cooperate or defect with their partner. The third domain is a three-player, public goods game where agents may choose to contribute all or none of their savings to a public pool which is then redistributed evenly with a growth multiplier of $1.3$ (compare Janssen & Ahn (2003)); payoff is measured in terms of player profits. In the fourth, we consider a three-player El Farol bar problem where players choose whether to go to a bar or stay home (Arthur, 1994). Agents receive maximum payoff ($2$) for attending an uncrowded bar (<3<3 people), followed by staying home ($1$), followed by attending a crowded bar (0). Lastly, we consider the classic Bach-Stravinsky game where agents must coordinate to attend a performance despite different preferences. We set $\gamma=0.99$ in all domains, and use the last joint action selected by each player as state, i.e., $\mathcal{S}=\mathcal{A}$.
5.1 Creativity
Our first application considers utilities that value solutions that cover more than a small subset of the state-action space, which leads to “creative” equilibria, as we show. We get such outcomes for utilities that get entropy bonuses for high (Shannon) entropy occupancy measures, $u(\mu)=r^{\top}\mu+\tau H(\mu)$.
To find equilibria of the original game, we anneal the weight on this entropy bonus towards zero, following prior work on homotopy methods for equilibria in Markov games (Eibelshäuser & Poensgen, 2019).
First, we consider a multiagent pathfinding problem. Two agents must coordinate to pass through a bottleneck doorway on their way to a joint goal state. The reward for reaching the goal state is $100$ for both players; 0.01-0.01 reward otherwise. Upon reaching the goal state, the agents are reset to the start state (leftmost grid in Figure 2). Our algorithm returned an approximate equilibrium where the final utility for each agent was $24.5$ and the exploitability was $1.7$ ($\approx 7\%$ of their utility).
A single rollout of the final learned policy is shown in Figure 2. The agents race to cross the doorway, after which, one agent takes the center position and the other steps aside. In the third and fourth to last frames, the blue agent moves downward due to the small remaining entropy bonus. Both agents then move upward towards the goal state. In the final frame, green executes a “no-op” action to the right as blue moves into goal position. Despite learning a factorized Nash equilibrium policy profile, the agents exhibit coordinated actions at certain step. This coordination is achieved through observations of partner players grid locations, but richer coordination is theoretically possible with richer observation spaces.
Next, we examine three iterated normal-form games. Figure 3 shows our algorithm has vanishing exploitability for all of them. In each game, directly minimizing exploitability in cvxpylayers crashes due to numerical instabilities. Round-robin descent exhibits a similar qualitative behavior to our method, and faster convergence than simultaneous descent.
For a study of creativity, it is valuable to inspect the equilibria the methods produced. Round-robin gradient descent converges to an asymmetric equilibrium in iterated El Farol where one player goes to the bar every night, while the two other players must alternate. In contrast, our approach reveals more nuanced, symmetric policies in IPD and PGG that we discuss below.
Note that measuring the entropy of a player’s occupancy measure is different from measuring the entropy of their policy; the latter only measures entropy of action distributions in each state, ignoring the distribution across states. Interestingly, this difference manifests in the structure of the equilibria we discover. At high entropy, agents must explore the entire state space, which includes joint cooperation in the iterated prisoner’s dilemma (IPD) and joint donation in the public goods game (PGG). As temperature is annealed, the player’s receive less of a bonus for exploration, however, this transient introduction to mutually beneficial play has a lasting impact on equilibrium selection.
In IPD, PGL finds a symmetric policy. In Table 2, we show player 0’s learned policy.
State $s=(a_{0}^{t-1},a_{1}^{t-1})$ $\operatorname*{arg\,max}_{a}\pi_{0}(a|s)$ $\max_{a}\pi_{0}(a|s)$
(Cooperate, Cooperate) Cooperate 0.73
(Cooperate, Defect) Defect 0.74
(Defect, Cooperate) Cooperate 0.66
(Defect, Defect) Defect 0.81
Table 2: (Approximate PGL Equilibrium on IPD) The utility per-player is $0.47$ compared to that of other classic IPD strategies: tit-for-tat ($0.5$), win-stay, lose-shift ($0.67$), grim trigger ($0.42$), defect-defect ($0.33$).
If both players cooperated on the last round, then they are likely to continue cooperating. If the other player defected, then you are likely to defect (even more likely if you defected rather than cooperated on the last round). If you defected and the other player cooperated, you are actually more likely to cooperate in the next round, in an act of reciprocation.
In the iterated public goods game, PGL finds a symmetric policy as well. In Table 3, we show player 0’s learned policy.
State $s=(a_{0}^{t-1},a_{1}^{t-1},a_{2}^{t-1})$ $\operatorname*{arg\,max}_{a}\pi_{0}(a|s)$ $\max_{a}\pi_{0}(a|s)$
(None, None, None) None 0.99
(None, None, All-In) None 0.66
(None, All-In, None) None 0.66
(None, All-In, All-In) All-In 0.60
(All-In, None, None) None 0.80
(All-In, None, All-In) All-In 0.56
(All-In, All-In, None) All-In 0.56
(All-In, All-In, All-In) All-In 0.86
Table 3: (Approximate PGL Equilibrium on iterated public goods game) Per-player utility is $0.03$ versus 0 for zero contribution policies.
In contrast to the zero contribution policies found by all other methods, we find a contribution policy that, intuitively, is more likely to contribute funds when other agents do.
We see that entropy bonuses in utilities allows us to select for creative equilibria.
5.2 Imitation
Our second application considers utilities that value policies similar to policies observed in human experiments (Romero & Rosokha, 2023, Table 1, Current, Direct-Response).
State $s=(a_{0}^{t-1},a_{1}^{t-1})$ $a^{*}=\operatorname*{arg\,max}_{a}\pi_{0}(a|s)$ $\pi_{0}(a^{*}|s)$ $\pi_{0}^{h}(a^{*}|s)$
(Cooperate, Cooperate) Cooperate 0.83 0.86
(Cooperate, Defect) Defect 0.52 0.65
(Defect, Cooperate) Defect 0.53 0.55
(Defect, Defect) Defect 0.86 0.87
Table 4: Approximate Nash equilibrium recovered by our algorithm in IPD after annealing KL regularization to the human policy reported in the rightmost column (Romero & Rosokha, 2023, Table 1, Current, Direct-Response). The utility per-player under our learned symmetric policy profile is $0.48$ versus $0.46$ for the human policy. In addition, our learned policy profile is $1.4\times 10^{-4}$-exploitable at every state, whereas the human policy profile is substantially more exploitable, being $0.47$-exploitable over any initial state. (Exploitability for all initial states may be seen as an analogue of Markov perfection in Markov games (Maskin & Tirole, 2001).
In this experiment, we build on the homotopy experiment in the creativity section where we annealed our entropy coefficient $\tau$, but instead of annealing entropy, we anneal a KL penalty to the human state-action occupancy measure, $u_{i}(\mu)=r^{\top}\mu-\tau d_{\operatorname{KL}}(\mu\|\mu_{\operatorname{ref}})$. Note that matching occupancy measures is akin to matching long-run trajectories whereas matching policies does not necessarily match trajectories when other agents adjust their policies as will be the case in our experiments over learning. The human policies were derived from experiments where subjects played the iterated prisoner’s dilemma, selecting cooperate (C) or defect (D) in each period (Romero & Rosokha, 2023, Table 1, Current, Direct-Response). Subjects were required to confirm their opponent’s action after each period. This ensured that they were capable of representing a policy that conditions on the previous action. Table 4 reports the symmetric policy we learned while regularizing to the human occupancy measure. Note that this new policy profile has slightly higher utility for all agents and is very close to an equilibrium, independent of the starting state.
After two scenarios where we used a sequence of cMGs to discover creative resp. human-like equilibria, we consider a setting in which the utilities are not annealed, but are constant across time.
5.3 Fairness
Our final application considers utilities that value fair visitation of states. In the Bach-Stravinsky game, two players must choose whether to attend a performance by Bach or Stravinsky. If they don’t meet, they get zero reward. However, one player prefers Bach to Stravinsky ($3$ vs. $2$), whereas the other player prefers Stravinsky to Bach ($3$ vs. $2$). We incorporate a term into both player’s objectives that penalizes $10$ times the squared difference between attending Bach and attending Stravinsky to incentivize fair, equal attendance of the two shows, $u_{i}(\mu)=r^{\top}\mu-(\sum_{a\in\{\operatorname{S},\operatorname{B}\}}\mu_{i}((\operatorname{B},\operatorname{B}),a)-\sum_{a\in\{\operatorname{S},\operatorname{B}\}}\mu_{i}((\operatorname{S},\operatorname{S}),a))^{2}$. We can write the second term as $\mu^{\top}\Sigma\mu$ for an appropriately chosen positive semidefinite matrix $\Sigma$.
In order to avoid the trivial equilibrium of voting $50/50$, we initialize logits for both players’ policies with a standard normal. We set temperature $\tau$ to zero and then optimize with a learning rate of $0.1$ for $1000$ iterations.
In $10$ random trials, both players converge to the same approximate NE where they vote for their favored event $60\%$ of the time regardless of their actions on the previous day. The maximum exploitability over the $10$ trials is $2.5\times 10^{-5}$, and the max difference between $\sum_{a}\mu_{i}((\operatorname{B},\operatorname{B}),a)$ and $\sum_{a}\mu_{i}((\operatorname{S},\operatorname{S}),a)$ is $2.14\times 10^{-5}$, implying this is a “fair” behavioral profile by our fairness metric.
6 Related Work
Our work relates to the single-agent literature on convex Markov games and equilibrium selection in Markov games. We rely on techniques from the theory of abstract economies, loss minimization for equilibrium computation, and use homotopy methods as inspiration for our experiments. Finally, we unify approaches to creativity, imitation, and fairness from multi-agent learning.
Convex Markov Decision Processes
Markov decision processes (MDPs) are the predominant framework for modeling sequential decision making problems, especially in infinite-horizon settings (Puterman, 2014). The goal of a decision maker in an MDP is typically to maximize a $\gamma$-discounted sum of rewards earned throughout the sequential decision process. In the infinite-horizon setting, recent research has exploited an alternative, but equivalent view of maximizing the expected reward under the agent’s stationary state-action occupancy measure (Zhang et al., 2020)—the probability of being in a given state and taking a given action. This viewpoint reveals an optimization problem with a linear objective (maximize return) and linear constraints (valid occupancy measure); from this launchpad, research has generalized to convex objectives that incorporate, for example, the (neg)entropy of the occupancy measure in order to maximize exploration of the MDP (Zahavy et al., 2021) or maximize robustness (Grand-Clément & Petrik, 2022).
Markov Games.
When multiple agents’ decision making problems interact, the MDP framework is extended to a Markov game, also known as a stochastic game (Thuijsman, 1997; Littman, 1994). In the game setting, we seek equilibria, a notion of simultaneous optimality for all agents. Fink (1964) proved the existence of stationary (time-independent) $\gamma$-discounted Nash equilibria in $n$-player, general-sum stochastic games. A homotopy approach that traces the continuum of quantal response equilibria performs well at approximating Nash equilibria in the limit of zero temperature (Eibelshäuser & Poensgen, 2019). Other approaches are tailored for more restricted two-player, zero-sum settings (Daskalakis et al., 2020; Goktas et al., 2023b) or settings where agent incentives are a priori aligned such as Markov potential games (Leonardos et al., 2021). Prior work proved a negative result for value iteration based approaches stating it is not possible to derive a stationary equilibrium policy from Q-values in general Markov games (Zinkevich et al., 2005).
Techniques.
In our existence proof, we rely on the theory of abstract economies of Arrow & Debreu (1954). Results on abstract economies have only recently found use in machine learning applications (Goktas & Greenwald, 2022; Goktas et al., 2023a; Jordan et al., 2023). Our application combines policy and occupancy measure views with general results on abstract economies.
Our proposed loss function extends that designed in recent work for normal-form games (NFGs) (Gemp et al., 2024) to the convex Markov game setting. In their work, the focus was on constructing a loss amenable to unbiased estimation. In NFGs, the projection matrix $\Pi_{T\mathcal{M}_{i}(\mu_{-i})}=\Pi_{T\mathcal{M}_{i}}$ is a constant matrix independent of player strategies. This allows the construction of an unbiased estimator of their loss as long as they can obtain unbiased gradients of player’s utilities. Our loss applies to a more general class of games, but sacrifices unbiasedness.
Our applications anneal the temperature of an entropy bonus and/or a Kullback-Leibler divergence penalty. Such approaches relate to homotopy continuation-based approaches to equilibrium computation. McKelvey & Palfrey (1995) introduced quantal response equilibria along with a homotopy from infinite to zero temperature defining their limiting logit equilibrium. They extended this to the extensive-form game setting (McKelvey & Palfrey, 1998), at which point others provided an in-depth study of its unique equilibrium selection properties (Harsanyi et al., 1988). More recent work extended this approach to Markov games (Eibelshäuser & Poensgen, 2019). Our approach selects equilibria using a homotopy-inspired method in our applications to creativity and imitation.
Applications.
Our goals of creativity, imitation, and fairness are not new to multi-agent applications. For example, Bakhtin et al. (2022) used KL-regularization towards human play to recover strategically superior policies in Diplomacy. Zahavy et al. (2022; 2023) leveraged convex MDPs to discover more creative play in Chess. And Hughes et al. (2018) models inequity aversion in complex multi-agent reinforcement learning domains. In contrast to these applications, our framework of convex Markov games allows for a unified analysis of several domains in a common language and using common algorithmic principles.
7 Conclusion
Convex Markov Games are a versatile framework for multi-agent reinforcement learning, in particular for equilibria that imitate human behavior, pursue creative equilibria, and those where additional constraints over states are desired. Not only do pure-strategy equilibria exist, there is also a differentiable upper bound for its optimization. Conceptually, optimization can be viewed via a parameterization as of policies and agent occupancy measures. The framework finds equilibria with diverse state visitation, that are close to human data, and that satisfy fairness constraints.
Convex Markov games open up several directions for future work, in particular in relation to decentralized training and settings where the transition dynamics are unknown. If dynamics are estimated, we pointed out issues obtaining unbiased estimates of the projection operator $\Pi_{T\mathcal{M}_{i}(\mu_{-i})}$. Our current approach computes an equilibrium for all agents using a centralized gradient. It seems believable that, given the fact that all agents solve convex optimization problems, decentralized training towards a notion of (coarse)-correlated equilibria in cMGs is possible.
Appendix A Equilibrium Proof
We now prove that a pure-strategy policy Nash equilibrium exists for any player utility functions that admit unique solutions in the space of occupancy measures (a convex, compact set). This includes strictly convex functions, but also the more general class of invex functions which may be non-convex.
For convenience, we repeat the Kakutani fixed point theorem here.
Theorem ((Kakutani, 1941)).
Let $\Pi=\prod_{i}\Pi_{i}$ be a non-empty, compact and convex subset of $\mathbb{R}^{|\mathcal{S}^{n}|\prod_{i}|\mathcal{A}_{i}|}$.
Let $\phi:\Pi\rightarrow 2^{\Pi}$ be a set-valued function on $\Pi$ with the following properties:
  • $\phi$ has a closed graph;
  • $\phi(\pi)$ is non-empty and convex for all $\pi\in\Pi$.
Then $\phi$ has a fixed point.
We now state a useful Lemma that proves every occupancy measure maps to a convex set of policies.
Lemma 2.
The set of policies that give rise to a single occupancy measure $\mu_{i}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}_{i}|}$ for a single player $i$ is a closed, convex set.
Proof.
The mapping from occupancy measures to policies in convex MDP settings is typically defined such that it is one-to-one even on states with zero occupancy measure, e.g.,
$\pi_{i}(\mu_{i})(a|s)$ $=\begin{cases}\frac{\mu_{i}(s,a)}{\sum_{a^{\prime}}\mu_{i}(s,a^{\prime})}&\text{ if }\sum_{a^{\prime}}\mu_{i}(s,a^{\prime})>0\\ 1/|\mathcal{A}_{i}|&\text{ else}.\end{cases}$ (10)
However, this definition is inadequate in a multi-player setting. In a game, we cannot assume (or force) a player’s policy takes a specific value unless it is rationalizable (i.e., they chosen that policy because it is their unqiue best response). Therefore, in the case where a state has zero occupancy, the set of allowable policies in that state is the entire simplex $\Delta^{\mathcal{A}_{i}}$; this is why we present equation 4 in Section 2.
In states where the occupancy has positive measure, the mapping from occupancy to policy is unique and results in a singleton, a closed, convex set. Otherwise, the mapping is to the simplex, also a closed, convex set. A player’s full policy is formed from the product sets of its state policies; the product of these convex sets is convex completing the claim. ∎
Corollary 1 (Convex Best Response Correspondence).
If a player’s best response correspondence is unique (a singleton) in the space of occupancy measures, then its best response correspondence in policy space is a convex set.
In what follows, although the theory of $n$-player, concave games by Rosen (1965) does not apply, we are able to trace a similar line of reasoning and employ Lemma 2 to prove existence of pure Nash equilibria. First, we recall some useful properties and present some preliminary definitions.
For $\pi\in\Pi$, $u_{i}(\pi)$ is continuous in $\pi$ from Lemma 3 ($\mu_{i}(\pi)$ is continuous) and continuity of the objective $u_{i}(\mu_{i})$.
An equilibrium point of the cMG is given by a point $\pi^{*}\in\Pi$ such that
$u_{i}(\pi^{*})$ $=\max_{\pi^{\prime}_{i}}\{u_{i}(\pi^{\prime}_{i},\pi^{*}_{-i})|(\pi^{\prime}_{i},\pi^{*}_{-i})\in\Pi\}(i=1,\ldots,n).$ (11)
The results to follow make use of the function $\rho(\pi,\pi^{\prime})$ defined for $(\pi,\pi^{\prime})\in\Pi\times\Pi$ by
$\rho(\pi,\pi^{\prime})$ $=\sum_{i=1}^{n}u_{i}(\pi^{\prime}_{i},\pi_{-i}).$ (12)
By construction, the maxima of $\rho$ in $\pi^{\prime}$ are equal to the product set of the maxima of each $u_{i}$ w.r.t. $\pi_{i}^{\prime}$ individually. We observe that for $(\pi,\pi^{\prime})\in\Pi\times\Pi$ we have $((\pi^{\prime}_{i},\pi_{-i}))\in\Pi$, $i=1,\ldots,n$, so that $\rho(\pi,\pi^{\prime})$ is continuous in $\pi$ and $\pi^{\prime}$ and by Corollary 1, its maximizers in $\pi^{\prime}$ form a closed, convex set for every fixed $\pi$, for $(\pi,\pi^{\prime})\in\Pi\times\Pi$. We now state our general theorem here.
Theorem 1 (Follows (Rosen, 1965)).
Given a convex Markov game with unique best response correspondences in the space of occupancy measures, Nash equilibria exist in stationary stochastic policies (and hence state-action occupancy measures).
Proof.
Consider the point-to-set mapping $\phi:\Pi\rightarrow\Pi$, given by
$\phi(\pi)$ $=\{\pi^{\prime}|\rho(\pi,\pi^{\prime})=\max_{z\in\Pi}\rho(\pi,z)\}.$ (13)
It follows from the continuity of $\rho(\pi,z)$ and the convexity of $\operatorname*{arg\,max}_{z\in\Pi}\rho(\pi,z)$ for fixed $\pi$ that $\phi$ is an upper semicontinuous mapping that maps each point of the convex, compact set $\Pi$ into a closed, convex subset of $\Pi$. Then by the Kakutani fixed point theorem, there exists a point $\pi^{*}\in\Pi$ such that $\pi^{*}\in\phi(\pi^{*})$, or
$\rho(\pi^{*},\pi^{*})$ $=\max_{z\in\Pi}\rho(\pi^{*},z).$ (14)
The fixed point $\pi^{*}$ is an equilibrium point satisfying equation 11. Suppose it was not. Then, for $i=\ell$, there would be a point $\pi_{\ell}=\bar{\pi}_{\ell}$ such that $\bar{\pi}=(\pi^{*}_{1},\ldots,\bar{\pi}_{\ell},\ldots,\pi^{*}_{n})\in\Pi$ and $u_{\ell}(\bar{\pi})>u_{\ell}(\pi^{*})$. But then $\rho(\pi^{*},\bar{\pi})>\rho(\pi^{*},\pi^{*})$, which contradicts equation 14. ∎
Appendix B Occupancy from Policy is Differentiable
Lemma 3.
Player $i$’s state-action occupancy measure $\mu_{i}$ is a differentiable (and hence continuous) function of the player policies $\pi$.
Proof.
Recall
$\mu_{i}(\pi)$ $=(1-\gamma)\Big([I-\gamma P^{\pi}]^{-1}\mu_{0}\mathbf{1}_{a_{i}}^{\top}\Big)\odot\pi_{i}$ (15)
where
$P^{\pi}(s^{\prime},s)$ $=\langle P_{j}^{\pi_{-j}}(s^{\prime},s,:),\pi_{j}(s,:)\rangle$ (16)
and
$P_{j}^{\pi_{-j}}(s^{\prime},s,:)$ $=\sum_{a_{-j}}P(s^{\prime}|s,a_{i},a_{-j})\prod_{k\neq j}\pi_{k}(s,a_{k})$ (17)
so
$P^{\pi}(s^{\prime},s)$ $=\sum_{\bm{a}}P(s^{\prime}|s,\bm{a})\prod_{j}\pi_{j}(s,a_{j}).$ (18)
Then,
$\frac{\partial\mu_{i}(x,y)}{\partial\pi_{j}(x^{\prime},y^{\prime})}$ $=(1-\gamma)\Big[[I-\gamma P^{\pi}]^{-1}\mu_{0}(s)\Big]_{x}\mathds{1}(x=x^{\prime},y=y^{\prime},j=i)$
$+(1-\gamma)\frac{\partial}{\partial\pi_{j}(x^{\prime},y^{\prime})}\Big([I-\gamma P^{\pi}]^{-1}\Big)(\mu_{0}\mathbf{1}_{a_{i}}^{\top})\odot\pi_{i}$ (19)
where
$\frac{\partial}{\partial\pi_{j}(x^{\prime},y^{\prime})}\Big([I-\gamma P^{\pi}]^{-1}\Big)$ $=\gamma[I-\gamma P^{\pi}]^{-1}\frac{\partial P^{\pi}}{\partial\pi_{j}(x^{\prime},y^{\prime})}[I-\gamma P^{\pi}]^{-1}$ (20)
and
$\frac{\partial P^{\pi}}{\partial\pi_{j}(x^{\prime},y^{\prime})}$ $=\begin{cases}0&\text{ if }x\neq x^{\prime}\\ P_{j}^{\pi_{-j}}(s^{\prime},x,y)&\text{ else}.\end{cases}$ (21)
Clearly, this requires inverting the matrix $I-\gamma P^{\pi}]$. Note that $P^{\pi}$ is a square state transition matrix with distributions on columns. By the same argument as Lemma 1, this matrix has full-row rank, and since it is square, it is non-singular, and hence invertible. Therefore, the derivative (19) always exists. ∎
Appendix C Continuity of Action Correspondence
It is known that the Bellman-flow constraints in Program (3) additionally encode a simplex constraint on $\mu_{i}$ (see this by summing the constraints over $s$). Note that the Bellman flow constraints in (3) encode a hyperplane. Therefore the feasible set is the intersection of a hyperplane with the simplex, resulting in a convex polytope. It is known that there always exists a stationary distribution $\mu_{i}$ satisfying these constraints. Note that the hyperplane has fixed dimension ($d_{h}=|\mathcal{S}||\mathcal{A}_{i}|-|\mathcal{S}|$) because the constraint matrix has full row-rank $|\mathcal{S}|$ (proven in Lemma 1 below).
Recall player $i$’s transition matrix given fixed policies $\pi_{-i}$ is:
$P_{i}^{\pi_{-i}}(s^{\prime}|s,a_{i})$ $=\sum_{a_{-i}\in\mathcal{A}_{-i}}P(s^{\prime}|s,a_{i},a_{-i})\prod_{j\neq i}\pi_{j}(a_{j}|s)$ (22)
where $a_{j}$ is player $j$’s action component of $a_{-i}$. Note this function is polynomial in $\pi_{-i}$ and therefore continuous in $\pi_{-i}$.
Note that the constraint matrix depends on the transition operator $P_{i}^{\pi_{-i}}$. Therefore, the orientation and translation of the hyperplane depend continuously on the entries in $P_{i}^{\pi_{-i}}$. As the hyperplane continuously moves, the boundary of the convex polytope induced by its intersection with the simplex also moves continuously.
In order for the policies to depend continuously on the other player occupancy measures, we need to ensure that the occupancies always have full support on all states. Recall that $\pi_{i}(a|s)=\frac{\mu_{i}(s,a)}{\sum_{a}\mu_{i}(s,a)}$ if $\sum_{a}\mu_{i}(s,a)>0$ else $1/|\mathcal{A}_{i}|$. For state-actions converging to zero occupancy measure, this would induce discontinuous changes in their induced policies which would induce discontinuities in the transition kernel for other agents. We ensure occupancies have full support by assuming $\mu_{0}(s)$ has full support.
Given the above intuition, we now rigorously prove continuity of the action correspondence, beginning with proving the Bellman flow constraints are all linearly independent (which guarantees its associated hyperplane has fixed dimension).
Lemma 1.
The Bellman flow constraint matrix has full row-rank ($|\mathcal{S}|$) and is fixed independent of other player policies.
Proof.
Note the Bellman flow constraints can be written in matrix form as
$\sum_{a\in\mathcal{A}_{i}}(I_{s^{\prime}\times s}-\gamma P_{i,a})\mu_{i,a}=(1-\gamma)\mu_{0}$ (23)
where $P_{i}=P(s^{\prime}|s,a)$ denotes the $\mathcal{S}\times\mathcal{S}\times\mathcal{A}_{i}$ tensor of transition probabilities and $P_{i,a}=P_{i,a}(s^{\prime}|s)$ selects out a single action, leaving an $\mathcal{S}\times\mathcal{S}$ matrix.
This constraint can be written without the $\sum_{a\in\mathcal{A}_{i}}$ by constructing the rectangular block matrices
$I_{s^{\prime}\times(sa)}$ $=\begin{bmatrix}I_{s^{\prime}\times s}\ldots I_{s^{\prime}\times s}\end{bmatrix}$ (24)
and
$P_{i,s^{\prime}\times(sa)}$ $=\begin{bmatrix}P_{i,a_{1}}\ldots P_{i,a_{m}}\end{bmatrix}$. (25)
Then
$(I_{s^{\prime}\times(sa)}-\gamma P_{i,s^{\prime}\times(sa)})\mu_{i}$ $=(1-\gamma)\mu_{0}$ (26)
We can examine the first $s^{\prime}\times s$ block of $(I_{s^{\prime}\times(sa)}-\gamma P_{i,s^{\prime}\times(sa)})$ and show that this matrix is full rank, i.e., of rank $|\mathcal{S}|$. If this matrix is full rank, then its rows are linearly independent. Extending our view to the full matrix, i.e., all columns, cannot render any of these original rows linearly dependent.
Note that the first block is represented by $(I_{s^{\prime}\times s}-\gamma P_{i,a})$ for some action $a$. Using Gershgorin’s circle theorem, we can bound the eigenvalues of this matrix to lie in a union of circles which all exclude the origin. Consider any column $c$, then every circle has a center in $\mathbb{R}_{+}$. In addition, the leftmost point of every circle lies in $\mathbb{R}_{+}$:
$(1-\gamma P_{i,a}(c|c)-\sum_{s^{\prime}\neq c}|\gamma P_{i,a}(s^{\prime}|c)|$ (27)
$=(1-\gamma P_{i,a}(c|c)-\gamma\sum_{s^{\prime}\neq c}P_{i,a}(s^{\prime}|c)$ (28)
$=1-\gamma\sum_{s^{\prime}}P_{i,a}(s^{\prime}|c)$ (29)
$=1-\gamma>0.$ (30)
Therefore, this matrix is non-singular, i.e., full-rank. There are only $|\mathcal{S}|$ rows, hence the row-rank cannot increase, which proves the claim. ∎
Lemma 4.
The action correspondence for each player is a continuous function of the other player’s policies.
Proof.
For any sequence of other player policies $\{\pi_{-i}^{(t)}\}_{t\in\mathbb{N}}$ converging to $\pi_{-i}$ and any $\mu_{i}$ in the feasible set induced by $\pi_{-i}$, define a corresponding sequence of player $i$ occupancy measures $\{\mu_{i}^{(t)}\}_{t\in\mathbb{N}}$ as follows.
Define a continuous function $\pi_{-i}(s):[1,\infty)\rightarrow\Pi_{-i}$ that linearly interpolates between each pair of subsequent elements in $\{\pi_{-i}^{(t)}\}_{t\in\mathbb{N}}$, e.g., $\pi_{-i}(s)=(1-(s-\lfloor s\rfloor))\pi_{-i}^{(\lfloor s\rfloor)}+(s-\lfloor s\rfloor)\pi_{-i}^{(\lceil s\rceil)}$ where $\lfloor\cdot\rfloor$ and $\lceil\cdot\rceil$ represent the rounding operations to the nearest integer (down/up respectively).
This continuous function $\pi_{-i}(s)$ induces a corresponding continuous function $P_{i}^{\pi_{-i}}(s)$ (see Section 2), defining the hyperplane that intersects the positive orthant at every $s$.
Note, we are guaranteed there always exists a feasible $\mu_{i}$ by the equivalence between the policy and occupancy measure views of MDPs.
Choose a basis $\{v_{j}(s)\}_{j\in[d_{h}]}$ with $v_{j}(s):[1,\infty)\rightarrow\mathbb{R}^{|\mathcal{S}||\mathcal{A}_{i}|}$ for the hyperplane that moves with the hyperplane (i.e., a fixed local reference frame for the hyperplane); select it such that $\mu_{i}$ lies in the direction of the limit one of these basis vectors $v_{k}^{\infty}=\lim_{s\rightarrow\infty}v_{k}(s)$ starting from the centroid $\mu_{i}^{c}$ of the feasible polytope induced by $\pi_{-i}$. Note that by Lemma 1 the dimensionality of this hyperplane $d_{h}$ is fixed, so it is clear how to construct such a basis.
Let $\mu_{i}^{c}(s):[1,\infty)\rightarrow\mathbb{R}^{|\mathcal{S}||\mathcal{A}_{i}|}$ be the centroid of the feasible polytope for all $s$. Define $b_{k}(s)$ to be the distance from $\mu_{i}^{c}(s)$ to the boundary in direction $v_{k}^{\infty}$ for all $s$. Note the hyperplane moves continuously, intersecting with the positive orthant, a convex set. This implies the feasible polytope remains convex for all $s$ with a continuously deforming boundary. In addition, we are guaranteed the feasible set is never empty by the equivalence between the policy and occupancy measure views of MDPs. Therefore, both $\mu_{i}^{c}(s)$ and $b_{k}(s)$ are continuous. Let $\mu_{i}^{c}=\lim_{s\rightarrow\infty}\mu_{i}^{c}(s)$, and similarly $b_{k}=\lim_{s\rightarrow\infty}b_{k}(s)$.
By this construction, $\mu_{i}$ can be represented as $\mu_{i}=\mu_{i}^{c}+\xi b_{k}v_{k}^{\infty}$ with $\xi\in[0,1]$.
Finally, define the sequence $\{\mu_{i}^{(t)}\}_{t\in\mathbb{N}}$ with each $\mu_{i}^{(t)}=\mu_{i}^{c}(t)(1/t)+(1-1/t)(\mu_{i}^{c}(t)+\xi b_{k}(t)v_{k}(t))$. This sequence always lies in the feasible set because it represents an interpolation between the centroid and the polytope boundary. Also, the sequence $\{\mu_{i}^{(t)}\}_{t\in\mathbb{N}}$ is built from an underlying set of continuous components; therefore, it converges to $\mu_{i}$ in the limit by construction: $\lim_{t\rightarrow\infty}\mu_{i}^{(t)}=\lim_{s\rightarrow\infty}\mu_{i}^{c}(s)(1/s)+(1-1/s)(\mu_{i}^{c}(s)+\xi b_{k}(s)v_{k}(s))=\mu_{i}^{c}+\xi b_{k}v_{k}^{\infty}=\mu_{i}$.
Lemma 5.
Assume that the initial state distribution $\mu_{0}$ has full support. Then, $\mathcal{M}_{i}:\times_{j\neq i}\Delta^{|\mathcal{S}|\times|\mathcal{A}_{j}|-1}\rightarrow\Delta^{|\mathcal{S}|\times|\mathcal{A}_{i}|-1}$ for each player $i$ is a continuous correspondence.
Proof.
Lemma 4 proves the action correspondence is continuous in the other players’ policies. Given every player’s state-occupancy has full-support, each player’s policy is a continuous function of their occupancy measure (see equation (4)). Therefore, the action correspondence is continuous over other players’ occupancy measures as well. ∎
Appendix D KKT Conditions Imply Fixed Point Sufficiency
Consider the following constrained optimization problem:
$\max_{\bm{x}\in\mathbb{R}^{d}}$ $\,\,f(\bm{x})$ (31a)
$s.t.\,\,g_{i}(\bm{x})$ $\leq 0\,\,\forall i$ (31b)
$h_{j}(\bm{x})$ $=0\,\,\forall j$ (31c)
where $f$ is concave and $g_{i}$ and $h_{j}$ represent inequality and equality constraints respectively. If $g_{i}$ and $h_{i}$ are affine functions, then any maximizer $\bm{x}^{*}$ of $f$ must satisfy the following necessary and sufficient KKT conditions (Ghojogh et al., 2021; Boyd & Vandenberghe, 2004):
  • Stationarity: $\mathbf{0}\in\partial f(\bm{x}^{*})-\sum_{j}\lambda_{j}\partial h_{j}(\bm{x}^{*})-\sum_{i}\mu_{i}\partial g_{i}(\bm{x}^{*})$
  • Primal feasibility: $h_{j}(\bm{x}^{*})=0$ for all $j$ and $g_{i}(\bm{x}^{*})\leq 0$ for all $i$
  • Dual feasibility: $\mu_{i}\geq 0$ for all $i$
  • Complementary slackness: $\mu_{i}g_{i}(\bm{x}^{*})=0$ for all $i$.
Lemma 6.
Assuming player $k$’s utility, $u_{k}(x_{k},x_{-k})$, is concave in its own strategy $x_{k}$, a strictly-positive primal-feasible strategy is a best response $\texttt{BR}_{k}$ if and only if it has zero projected-gradient norm.
Proof.
Consider the problem of formally computing $\epsilon_{k}(\bm{x})=\max_{z\geq 0,Az=b}u_{k}(z,x_{-k})-u_{k}(x_{k},x_{-k})$:
$\max_{z\in\mathbb{R}^{d}}$ $\,\,u_{k}(z,x_{-k})-u_{k}(x_{k},x_{-k})$ (32a)
$s.t.-z_{k}$ $\leq 0\,\,\forall k$ (32b)
$A_{j}z-b_{j}$ $=0\,\,\forall j.$ (32c)
Note that the objective is linear (concave) in $z$ and the constraints are affine, therefore the KKT conditions are necessary and sufficient for optimality. Recall that we assume that the solution $z^{*}$ is positive, $z^{*}_{k}>0$ for each $k$. Also, let $e_{k}$ be a onehot vector, i.e., a zeros vector except with a $1$ at index $k$. Mapping the KKT conditions onto this problem yields the following:
  • Stationarity: $\mathbf{0}\in\partial u_{k}(z^{*},x_{-k})-\sum_{j}\lambda_{j}A_{j}+\sum_{k}\mu_{k}e_{k}$
  • Primal feasibility: $A_{j}z=b_{j}$ for all $j$
  • Dual feasibility: $\mu_{i}\geq 0$ for all $k$
  • Complementary slackness: $\mu_{k}z^{*}_{k}=0$ for all $k$
where $\partial u_{k}(z,\cdot)$ is the subdifferential at $z$. Consider any primal-feasible point $Az^{*}=b$. Given our assumption that $z^{*}_{k}>0$, by complementary slackness and dual feasibility, each $\mu_{k}$ must be identically zero. This implies the stationarity condition can be simplified to $\mathbf{0}\in\partial u_{k}(z^{*},x_{-k})-\sum_{j}\lambda_{j}A_{j}=\partial u_{k}(z^{*},x_{-k})-A^{\top}\lambda$. Rearranging terms we find that for any $z^{*}$, there exists $\lambda$ such that
$\sum_{j}\lambda_{j}A_{j}$ $\in\partial u_{k}(z^{*},x_{-k}).$ (33)
Equivalently, elements of $\partial u_{k}(z^{*},x_{-k})$ are in the row-span of $A$.
For the rest of the proof, we follow the derivation of the gradient projection method (Luenberger et al., 1984, Sec 12.4, p 364). Let $\nabla_{k}\in\partial u_{k}$ be a subderivative (gradient), i.e., an element of the subdifferential.
Note that, in general, we can write any gradient as a sum of elements from the row-span of $A$ and its orthogonal complement $d_{k}$, which lies in the tangent space of the feasible set:
$\nabla_{k}=d_{k}+A^{\top}\lambda.$ (34)
We may solve for $\lambda$ through the requirement that $Ad_{k}=0$, i.e., any movement within the tangent space of the feasible set remains feasible. Thus
$Ad_{k}$ $=A\nabla_{k}-(AA^{\top})\lambda=0$ (35)
$\implies\lambda$ $=(AA^{\top})^{-1}A\nabla_{k}$ (36)
$\implies d_{k}$ $=\nabla_{k}-A^{\top}\lambda=[I-A^{\top}(AA^{\top})^{-1}A]\nabla_{k}$ (37)
$=\Pi_{TA}(\nabla_{k})$ (38)
where $\Pi_{TA}[I-A^{\top}(AA^{\top})^{-1}A]$ is the matrix that projects any gradient vector onto the tangent space of the feasible set given by the constraint matrix $A$.
The fact that elements of $\partial u_{k}(z^{*},x_{-k})$ are in the row-span of $A$ implies that $0=d_{k}=\Pi_{TA}(\nabla_{k})$ necessarily, completing the claim. ∎
Appendix E Entropy Regularized Loss Bounds Exploitability
Our derived exploitability bound only requires concavity of the utility, bounded diameter of the feasible set, and linearity of feasible constraints (i.e., feasible set is subset of hyperplane). In what follow, let player $k$’s loss be the negative of their utility, i.e., $\ell_{k}=-u_{k}$.
Lemma 7.
The amount a player can gain by deviating is upper bounded by a quantity proportional to the norm of the projected-gradient:
[…]
Proof.
Let $z$ be any point in the feasible set. First note that $\Pi_{T\mathcal{M}_{k}}(z)=Bz$ where $B$ is an orthogonal projection matrix; this implies $B^{2}=B=B^{\top}$. Then by convexity of $\ell_{k}$ with respect to $z$,
$\ell_{k}(\bm{\mu})-\ell_{k}(z,\mu_{-k})$ $\leq(\nabla^{k}_{\mu_{k}})^{\top}(z-\mu_{k})$ (40a)
$=(\nabla^{k}_{\mu_{k}})^{\top}\Pi_{T\mathcal{M}_{k}}(\underbrace{z-\mu_{k}}_{\in T\mathcal{M}})$ (40b)
$=(\underbrace{\Pi_{T\mathcal{M}_{k}}(\nabla^{k}_{\mu_{k}})}_{Bz=B^{\top}z})^{\top}\underbrace{(z-\mu_{k})}_{\texttt{Diam}(\mathcal{M}_{k})\leq\sqrt{2}}$ (40c)
[…]
where the first equality follows from the fact any two points $z$ and $\mu_{k}$ lying in the same hyperplane, by definition, form a direction lying in the tangent space of the hyperplane. The second equality follows from symmetry of the projection operator and simply grouping its application to the left hand term; we also note that the feasible set is a subset of the simplex, which has a diameter of $\sqrt{2}$. Finally, the last step follows from Cauchy-Schwarz.
Theorem 2 (Low Temperature Approximate Equilibria are Approximate Nash Equilibria).
Let $\nabla^{k\tau}_{\mu_{k}}$ be player $k$’s entropy regularized gradient and $\bm{\mu}$ be an approximate equilibrium of the entropy regularized game. Then it holds that
[…]
Proof.
Beginning with the definition of exploitability, we find
$\ell_{k}(\bm{\mu})-\ell_{k}(\texttt{BR}_{k},\mu_{-k})$ $=\big(\ell_{k}(\bm{x})+\tau S(\mu_{k})-\tau S(\mu_{k})\big)$ (42a)
$\qquad-\big(\ell_{k}(\texttt{BR}_{k},\mu_{-k})+\tau S(\texttt{BR}_{k})-\tau S(\texttt{BR}_{k})\big)$
$=\ell_{k}^{\tau}(\bm{\mu})-\ell_{k}^{\tau}(\texttt{BR}_{k},\mu_{-k})+\tau\big(S(\mu_{k})-S(\texttt{BR}_{k})\big)$ (42b)
$\leq\ell_{k}^{\tau}(\bm{\mu})-\min_{z\in\mathcal{M}_{k}}\ell_{k}^{\tau}(z,\mu_{-k})+\tau\max_{z^{\prime}\in\mathcal{M}_{k}}S(z^{\prime})$ (42c)
[…]
where the second equality follows from the definition of player $k$’s entropy regularized loss $\ell_{k}^{\tau}$, the first inequality from nonnegativity of entropy $S$, the second inequality from convexity of $\ell_{k}^{\tau}$ with respect to its first argument (Lemma 7), and the last from the maximum possible value of Shannon entropy over distributions on $|\mathcal{S}||\mathcal{A}_{k}|$ elements. ∎
Corollary 2 ($\mathcal{L}^{\tau}$ Scores Nash Equilibria).
Let $\mathcal{L}^{\tau}(\bm{\mu})$ be our proposed entropy regularized loss function and $\bm{\mu}$ be any strategy profile. Then it holds that
$\epsilon$ $\leq\tau\log\Big(|\mathcal{S}|^{n}|\mathcal{A}|\Big)+\sqrt{2n\mathcal{L}^{\tau}(\bm{\mu}}).$ (43)
Proof.
Beginning with the definition of exploitability and applying Lemma 2, we find
$\epsilon$ $=\max_{k}\max_{z}\ell_{k}(\bm{\mu})-\ell_{k}(z,\mu_{-k})\qquad\text{(recall each $\epsilon_{k}\geq 0$)}$ (44a)
$\leq\sum_{k}\max_{z}\ell_{k}(\bm{\mu})-\ell_{k}(z,\mu_{-k})$ (44b)
[…]
$=\tau\log\Big(|\mathcal{S}|^{n}|\mathcal{A}|\Big)+\sqrt{2}\Big|\Big|||\Pi_{T\mathcal{M}}(\nabla^{1\tau}_{\mu_{1}})||_{2},\ldots,||\Pi_{T\mathcal{M}}(\nabla^{n\tau}_{\mu_{n}})||_{2}\Big|\Big|_{1}$ (44f)
$\leq\tau\log\Big(|\mathcal{S}|^{n}|\mathcal{A}|\Big)+\sqrt{2n}\Big|\Big|||\Pi_{T\mathcal{M}}(\nabla^{1\tau}_{\mu_{1}})||_{2},\ldots,||\Pi_{T\mathcal{M}}(\nabla^{n\tau}_{\mu_{n}})||_{2}\Big|\Big|_{2}$ (44g)
[…]
$=\tau\log\Big(|\mathcal{S}|^{n}|\mathcal{A}|\Big)+\sqrt{2n\mathcal{L}^{\tau}(\bm{\mu}}).$ (44i)
Appendix F Game Matrices
We provide the payoffs for the prisoner’s dilemma and Bach-Stravinsky game used in experiments.
     $C$         $D$
$C$ 1,1-1,-1 3,0-3,0
$D$ 0,30,-3 2,2-2,-2
$\implies$      $C$         $D$      $C$ $\nicefrac{{2}}{{3}},\nicefrac{{2}}{{3}}$ $0,1$ $D$ $1,0$ $\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}}$
Figure 4: Prisoner’s Dilemma Game: We shift and normalize the payoffs of the prisoner’s dilemma game (left) to be positive and with maximum value 11 (right).
     $C$         $D$
$C$ $3,2$ $0,0$
$D$ $0,0$ $2,3$
Figure 5: Bach-Stravinsky Game.
Appendix G Obstacles to a General cMG NE-Existence Proof
Table 5 summarizes the obstacles to an NE-existence proof for general concave utilities.
Theory N.C. of $u_{i}(\pi_{i})$ N.C. of BR N.C. of $\mathcal{U}$ $\mathcal{U}\neq(\operatorname*{\vartimes}_{i=1}^{n}\mathcal{U}_{i})$ Deviations
Kakutani NA $\times$ NA NA NA
Concave Games $\times$ NA $\times$ NA NA
Hidden Games NA NA NA $\times$ NA
Abstract Economy NA NA NA NA $\times$
Table 5: Summarization of obstacles to various Nash equilibrium existence proof techniques. N.C. stands for non-convexity or non-concavity where appropriate. An $\times$ means the theory fails because of the issue. “Deviations” means that a notion of reasonable player deviations in the MDP are not allowed by the theory.
G.1 Non-Convexity of Best Response Correspondences
We consider a simple $2$-player, 22-state, 22-action, symmetric convex Markov game. If both agents select action 0, they transition from their current state to the other state; otherwise, they remain put. We set $\gamma=0.95$ and the initial state measure $\mu_{0}$ to be uniform.
Let player $2$ employ one of the following nearby policies (I-III):
$\pi_{2}^{I}(a_{0}|s)$ $=\begin{cases}0.40&\text{if $s=s_{0}$}\\ 0.80&\text{else}\end{cases},$
$\pi_{2}^{II}(a_{0}|s)$ $=\begin{cases}0.45&\text{if $s=s_{0}$}\\ 0.80&\text{else}\end{cases},$
$\pi_{2}^{III}(a_{0}|s)$ $=\begin{cases}0.40&\text{if $s=s_{0}$}\\ 0.75&\text{else}\end{cases}$
where player $2$’s probabilities for action $a_{1}$ are implied.
In addition, consider the following “safe” MARL problem where we designate safe long-run visitation measure of certain states (e.g., Bach-Bach or Stravinsky-Stravinsky) and usage of certain actions (e.g., a safe range of motor torques); equivalently, we rule out certain unsafe states and actions. Note this framework could also be used to tackle the fairness application we explored in Section 5.3. As an example of “safe” MARL, define the following convex loss (negative utility) over occupancy measures for player $1$:
[…]
where $\mu_{a}$ and $\mu_{s}$ are player $1$’s action and state marginals respectively. The target measures $t_{a}$ and $t_{s}$ are used along with the radii $\nicefrac{{1}}{{20}}$ and $\nicefrac{{1}}{{4}}$ to encode the regions of “safe” occupancy measures. If player $1$ deviates from either region by more than the radius (as measured by the infinity norm), then they accrue a loss. Otherwise, player $1$’s loss is zero.
Figure 6 reveals the shapes of player $1$’s best response correspondences in this setting, showing in particular, that its image in policy space is non-convex for each of the policies. This feature of the problem breaks the assumptions of the Kakutani fixed point theorem rendering a Nash equilibrium existence proof non-trivial, yet desireable.
(a) Occupancy Measure Space (I)
(b) Policy Space (I)
(c) Occupancy Measure Space (II)
(d) Policy Space (II)
(e) Occupancy Measure Space (III)
(f) Policy Space (III)
Figure 6: Non-convexity of best response correspondences (black regions) in policy space under concave utilities for a “safe” MARL application. Roman numerals indicate the policy employed by player 22 in each plot.
G.2 Non-Convexity of Joint Strategy Set
We again consider the same convex Markov game as in Section G.1. Note that the feasible set of joint occupancy measures is entirely defined by the parameters of the transition kernel, discount factor, and initial state measure (reward is irrelevant to the Bellman flow constraints).
Figure 7: (The set of feasible occupancy measures is non-convex) To be able to visualize the feasible set, we consider symmetric occupancy measures (occupancy is the same for both players). We build a probability mesh ranging from $0.01$ probability to $0.99$ probability for all state-action occupancies $\mu(s,a)$, and plot all points where the norm of the constraint violation vector is below $0.01$. We then identify two points in this set and draw a line segment between them. The midpoint of this line segment is a convex combination of the two endpoints and should necessarily lie in the feasible set if it is convex. However, as one can clearly see from the figure, the midpoint lies well outside the feasible set and we can confirm its constraint violation norm is above $0.3$.
The set of joint occupancy measures is non-convex. This breaks the assumption made in the theory of $n$-player concave games (Rosen, 1965). It also breaks most known results on monotone quasi-variational inequalities (Kim, 2021; Facchinei & Kanzow, 2007).
G.3 Hidden Games
In hidden games (Sakos et al., 2024), players’ strategy spaces are independently mapped from some complex space to a latent strategy space where payoffs are directly calculated. In (Sakos et al., 2024, Definition 1), they assume there exists a map $\mu_{i}=\mu_{i}(\pi_{i})$ for our setting. Note, in cMGs, $\mu_{i}=\mu_{i}(\pi_{1},\ldots,\pi_{n})$, and so it fails this condition. In addition, they assume a latent space $\mathcal{U}=\operatorname*{\vartimes}_{i}\mathcal{U}_{i}$, i.e., the joint strategy set is a product space. This also fails in cMGs.
G.4 Abstract Economies
In an abstract economy (Arrow & Debreu, 1954), players’ strategy sets may depend on the actions chosen by other players; generally, each player is tasked with the problem:
$\max_{\mu_{i}\in\mathcal{M}^{\prime}_{i}(\mu_{-i})}u_{i}(\mu_{i},\mu_{-i})$
where $\mathcal{M}^{\prime}_{i}(\mu_{-i})$ maps every profile of opponent occupancy measures to a feasible set of occupancy measures for player $i$, the so-called action correspondence. Note that unlike $\mathcal{M}_{i}(\mu_{-i})$ defined in Section 3, $\mathcal{M}^{\prime}_{i}(\mu_{-i})$ must respect the occupancy measure constraints of all players. Intuitively, player $i$’s deviation to an occupancy measure $\mu_{i}^{\prime}$ cannot result in a modification to $\mu_{j}$ for any $j\neq i$. All players share the same state occupancy marginals so this means that player $i$ cannot change their long-run state visitation measure by deviating. This is overly restrictive as it means that player $i$ cannot increase their utility in many games of interest, for example in games with reward functions that depend only on state. For this reason, we consider deviations in occupancy measures, and in turn, the abstract economy model insufficient for cMGs.